A 10-year corporate bond with a total face value of $25,000,000 and a stated coupon rate of 7.9% APR
payable semi-annually was issued at a price to yield a return of 8.4% to investors. What was the
proceeds from the sale of the bond and the quoted bond price at issuance?

N:

I/Y:

PV:

PMT:

FV:

Mode:

Excel Formula:

Bond Proceeds:

Answer:
Answer #1

N = 10*2=20
I/Y = 8.4%/2=4.2%
PMT = 7.9%*25,000,000/2 = 987,500
FV = 25,000,000
EXCEL Formula = PV(4.2%,20,-987500,-25000000) = 24,165,451.04

Bond Proceeds = 24,165,451.04

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well